OpenStudy (anonymous):

The Jurassic Zoo charges $10 for each adult admission and $3 for each child. The total bill for the 145 people from a school trip was $743. How many adults attended and how many children?

OpenStudy (texaschic101):

a = adult c = child a + c = 145 --> c = 145 - a 10a + 3c = 743 now sub 145 - a in for c in the 2nd equation can you handle that ?

OpenStudy (anonymous):

im getting a decimal

OpenStudy (texaschic101):

let me check...hold on

OpenStudy (texaschic101):

10a + 3c = 743 10a + 3(145 - a) = 743 10a + 435 - 3a = 743 7a + 435 = 743 7a = 743 - 435 7a = 308 a = 308/7 a = 44 I did not get a decimal

OpenStudy (texaschic101):

now just sub 44 in for a in the first equation and solve for c

OpenStudy (anonymous):

ohh ok, i see where i went wrong. i had 10a+435+3a=743

OpenStudy (texaschic101):

oh..-3a not + 3a...

OpenStudy (texaschic101):

can you solve for c now ?

OpenStudy (texaschic101):

just sub 44 in for a in the first equation and solve for c

OpenStudy (texaschic101):

you can then check your answers by subbing in what you got for a and c into the second equation and if the answer comes out equal, you have done it correctly

OpenStudy (anonymous):

i got c=101

OpenStudy (texaschic101):

yes....now check your answers

OpenStudy (anonymous):

checks out. 44+101=145 :)

OpenStudy (texaschic101):

10a + 3c = 743 10(44) + 3(101) = 743 finish this and what do you get ?